Calculate the nuclear binding energy and nuclear binding energy per nucleon



1
Expert's answer
2021-03-25T04:09:34-0400

Binding Energy "(B.E) =\u2206m\u00d7c^2"

Where ‘∆m’ is the mass defect and ‘c' is the speed of light


Binding energy per nucleon "(B.E.N) = \frac{∆m×c^2}{A} = \frac{B.E}{A}"


Where 'A' is number of necleouns in nucleus